### v3

# Soft-TransFormer for Continual Learning

Abstract
Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformer (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer. Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention. These masks enable smooth and stable task adaptation while preserving shared representations. Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF). Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
Machine Learning, Continual Learning, Sparse TransFormers
1 Introduction
Continual Learning (CL), also known as Lifelong Learning (Thrun, 1994; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), aims to endow models with the ability to acquire new knowledge over time without erasing previously learned information. An ideal CL system would enable both forward transfer, where prior knowledge facilitates learning new tasks, and backward transfer, where new experiences refine earlier representations. However, modern neural networks remain highly susceptible to catastrophic forgetting (McCloskey and Cohen, 1989), where updates from new tasks interfere with parameters essential for earlier ones, resulting in significant performance degradation.
A large body of prior work - predominantly developed for convolutional neural networks - tackles catastrophic forgetting through three major strategies: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) constrain updates to preserve knowledge of previously learned tasks; (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) store or synthesize past data for replay; and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022) reduce inter-task interference via task-specific modules or structural expansions.
Figure 1: Soft-TransFormers (Soft-TF): the objective is to design a fully fine-tuned model that works well across multiple continual learning settings with incurring task-wise soft network training of attention and feedforward networks, leveraged by WLTH.
With the advent of large-scale pre-trained Transformers, i.e., ViT (Dosovitskiy et al., 2021), a new paradigm - (4) Prompt-based Continual Learning—has emerged. By freezing the backbone and learning lightweight prompts, these methods adapt the model to sequential tasks while preserving the strong representational priors of the pre-trained architecture. L2P (Wang et al., 2022d) first demonstrated that prompt retrieval can bridge prompting and continual representation learning. DualPrompt (Wang et al., 2022c) further decomposes prompts into task-general and task-specific components, improving adaptation and stability. Subsequent advances include DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), LGCL (Khan et al., 2023), ESN (Wang et al., 2022b), CPrompt (Gao et al., 2024), and NoRGa (Le et al., 2024).
Despite their effectiveness, prompt-based CL methods exhibit several important limitations. First, performance depends heavily on the quality, granularity, and design of prompts, making the method sensitive to prompt initialization and capacity. Second, as the number of tasks grows, managing a large and diverse prompt pool becomes cumbersome. Third, prompt tuning inherently lacks the expressiveness of full fine-tuning-restricted prompts alone cannot capture the fine-grained task-specific variations needed for diverse and uncorrelated sequential tasks. While frozen pre-trained models offer strong global priors, they do not directly provide task-adaptive pathways, limiting the flexibility of prompt-only tuning. These limitations motivate the exploration of alternative fine-tuning paradigms that combine the stability of frozen backbones with the flexibility of full parameter adaptation.
To address these challenges, we revisit the parameter-efficient fine-tuning landscape, especially in comparison to methods such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), as illustrated in Figure˜2. We focus on two key desiderata in sequentially adapting pre-trained foundation models: (1) minimizing catastrophic forgetting and (2) ensuring parameter-efficiency during continual adaptation. Motivated by these goals, we introduce the following hypothesis:
Well-initialized Lottery Ticket Hypothesis (WLTH). A well-initialized dense neural network contains structured, globally stable solutions that preserve prior task knowledge, while also exposing local subspaces that can be selectively adapted for new tasks. These subspaces - corresponding to soft, real-valued subnetworks - enable isolated and parameter-efficient fine-tuning without overwriting the information encoded in the frozen backbone.
Inspired by WLTH, we propose Soft-Transformer (Soft-TF), a new parameter-efficient continual learning framework that discovers task-specific soft subnetworks within well-trained Transformer backbones, as shown in Figure˜1. Soft-TF incrementally learns task-adaptive multiplicative masks that modulate key, query, value, and output projections in self-attention layers, enabling flexible and stable adaptation with minimal interference between tasks.
Our main contributions are as follows:
  • We propose Soft-Transformer (Soft-TF), a new continual learning paradigm inspired by the Well-initialized Lottery Ticket Hypothesis, which learns compact task-specific soft subnetworks from well pre-trained Transformers.
  • We demonstrate that Soft-TF achieves state-of-the-art parameter-efficient continual learning performance, outperforming DualPrompt, NoRGa, Adapter, and LoRA across class-incremental learning scenarios.
Refer to caption Refer to caption Refer to caption
(a) Adapter (b) LoRA (c) our Soft-TransFormer
Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2022) are trained at all attention layers. Note that all pre-trained parameters $\bm{\theta}$ are fixed in fine-tuning t-th task parameter $\bm{m}_{t}$ in Continual Learning (CL) scenarios.
2 Related Work
Continual Learning. Continual Learning (CL) aims to enable models to accumulate knowledge over a sequence of tasks while preserving performance on previously learned ones (McCloskey and Cohen, 1989; Thrun, 1994). Early efforts focused on mitigating catastrophic forgetting, the tendency of neural networks to override earlier representations when trained on new data. Regularization-based methods address this by constraining parameter updates that would negatively impact past tasks, often by estimating parameter importance or imposing functional stability (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020). Rehearsal-based approaches alleviate forgetting through explicit memory replay, either by storing exemplar sets or synthesizing pseudo-data from generative models (Rebuffi et al., 2017; Chaudhry et al., 2019a; Buzzega et al., 2020). Architecture-based strategies expand or isolate parts of the model so that new tasks use disjoint or sparsely overlapping parameter subsets, thereby preventing interference (Mallya et al., 2018; Serrà et al., 2018; Wortsman et al., 2020; Kang et al., 2022). Although these methods have proven effective, they often require additional storage, dynamic model growth, or explicit replay buffers, which can become impractical in long-horizon or resource-constrained CL scenarios.
Parameter-efficient Adaptation. The shift from training models from scratch toward adapting large-scale pre-trained Transformers - such as ViT and CLIP (Dosovitskiy et al., 2021; Radford et al., 2021) - has significantly influenced recent CL research. Parameter-efficient fine-tuning (PEFT) methods, including adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), introduce compact trainable modules while keeping most backbone parameters frozen. These approaches provide stronger adaptation capacity than traditional CL techniques but still accumulate per-task modules or suffer from interference across tasks when task similarity is low. Prompt-based CL offers an even lighter adaptation mechanism: instead of modifying model weights, prompts act as learnable soft instructions prepended or inserted into the Transformer architecture. L2P (Wang et al., 2022d) initiated this direction by retrieving task-relevant prompts from a pool, while DualPrompt (Wang et al., 2022c) decomposed prompts into task-general and task-specific components to balance stability and plasticity. More recent advances aim to improve the expressiveness and robustness of prompt-based continual learning. CODA-Prompt (Smith et al., 2023b) utilizes decomposed attention to assemble input-conditioned prompts, mitigating task interference, while S-Prompt (Wang et al., 2022a) emphasizes simplicity and stability via domain-independent prompts. HiDe-Prompt (Wang et al., 2023) explores hierarchical and disentangled prompt representations to better capture task structures, and NoRGa (Le et al., 2024) introduces non-linear residual gating to enhance adaptability by connecting prompting with mixture-of-experts mechanisms. Collectively, these methods highlight the potential of prompt-based PEFT for continual learning, while also exposing inherent limitations in relying solely on prompt-level conditioning for scalable and expressive task adaptation.
Limitations of Prompt-based CL. Despite their attractive efficiency, prompt-based methods face intrinsic constraints that limit their robustness in practical continual learning settings. Since prompts modify representations only implicitly through input conditioning, their expressiveness is fundamentally lower than module-level or weight-level adaptation, making them less suited for highly heterogeneous or weakly correlated task sequences. Moreover, prompt retrieval and initialization are sensitive to distribution shifts, and prompt pools may grow without bound as tasks accumulate, reducing scalability and increasing memory cost. Adapter- and LoRA-based PEFT methods provide stronger adaptation capacity but may introduce task interference or undesirable parameter growth. These challenges highlight the need for a principled and expressive mechanism that maintains the efficiency of PEFT while allowing direct modulation of the pre-trained backbone. Motivated by the Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a Soft-Transformer (Soft-TF), which learns soft, real-valued subnetworks that multiplicatively modulate attention projections in a frozen Transformer. By selectively activating task-relevant subspaces within a well-initialized model, Soft-TF achieves stable, expressive, and parameter-efficient continual adaptation without relying solely on prompt proliferation.
Figure 3: Soft-TransFormer (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-MSAs are trained together with a classifier. At test time, an input is transformed by a query function (Prompt ID) or task identifier to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$.
3 Prerequisites
We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2021) in Class Incremental Learning (CIL) scenarios.
3.1 Problem Statement
Continual Learning (CL) involves training deep neural networks (DNN) on time-variant data represented as a sequence of tasks, $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{\mathcal{T}}\}$. Each $t$-th task, $\mathcal{D}_{t}=\{(\bm{x}_{i}^{t},y_{i}^{t})^{n_{t}}_{i=1}\}$ consists of $n_{t}$ tuples where $\bm{x}_{i}^{t}\in\mathcal{X}_{t}$ is an input sample and $y_{i}^{t}\in\mathcal{Y}_{t}$ is the corresponding label. When a task $\mathcal{X}_{t}$ arrives, a model $f_{\theta}$ is trained for the current task, while data from previous tasks is inaccessible. This work focuses primarily on class incremental learning (CIL), in which the task-ID is not given during inference.
3.2 Prompt-based Class Incremental Learning (CIL)
A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022d) is first proposed. In this model, a prompt $\bm{p}$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting. To select suitable prompts for task-specific training, L2P utilizes a prompt pool $\bm{P}$ containing numerous prompt-key pairs, $\{\bm{p}_{t},\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$, where $\bm{p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $\bm{k}_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs. Building on L2P, DualPrompt (Wang et al., 2022c) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning. DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in CIL scenarios.
4 Soft-TransFormer
This section presents Soft-TransFormer (Soft-TF), a parameter-efficient continual learning framework that adapts a frozen pre-trained Transformer through continuous, learnable soft-subnetworks and structured multi-layer prompting. Soft-TF preserves the rich representational structure encoded in the pre-trained backbone while enabling flexible task-specific modulation with minimal computational overhead and significantly reduced forgetting.
4.1 Soft-MSA Layers
Let $\bm{\theta}$ denote the parameters of a pre-trained Transformer, which remain fixed throughout continual learning. For each task $t$, as depicted in Figure˜3, Soft-TF introduces a set of real-valued masks
$\bm{m}_{t}=\{\bm{m}_{t}^{Q},\,\bm{m}_{t}^{K},\,\bm{m}_{t}^{V},\,\bm{m}_{t}^{O}\},$
which multiplicatively modulate the attention projections. Given hidden features $h$, we obtain:
$\bm{Q}=h(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q}),\;\bm{K}=h(\bm{\theta}^{K}\odot\bm{m}_{t}^{K}),\;\bm{V}=h(\bm{\theta}^{V}\odot\bm{m}_{t}^{V}).$
The output of the multihead attention layer is:
$\mathrm{MSA}(h)=\mathrm{Attn}(\bm{Q}_{i},\bm{K}_{i},\bm{V}_{i})\Big(\bm{\theta}^{O}\odot\bm{m}_{t}^{O}),$ (1)
where $\mathrm{Attention}(\cdot)$ denotes standard scaled dot-product attention.
Intuition. The masks are initialized at $\mathbf{1}$, ensuring that the initial computation matches the pre-trained model exactly. During optimization, the masks induce task-specific, continuous subnetwork selection within each attention projection. This allows Soft-TF to adjust representational pathways while keeping the global structure of the backbone fixed, thereby constraining parameter drift and enabling stable adaptation across tasks.
4.2 Prompts Coupled with Soft-Subnetwork
Soft-TF incorporates both general prompts (shared across tasks) and expert prompts (task-specific) into a multi-layer prompting mechanism. Let $\bm{g}=\{\bm{g}^{l}\}_{l=l_{g}^{\text{start}}}^{l_{g}^{\text{end}}}$, $\bm{e}_{t}=\{\bm{e}_{t}^{l}\}_{l=l_{e}^{\text{start}}}^{l_{e}^{\text{end}}}$ denote the layerwise G-prompts and E-prompts, respectively, with the constraint $l_{g}^{\text{start}},l_{g}^{\text{end}}]\cap[l_{e}^{\text{start}},l_{e}^{\text{end}}]=\emptyset.$
Given an input sequence $\bm{x}$, prompts are concatenated as
$\bm{z}=[\,\bm{p};\,\bm{x}\,],\quad\bm{p}\in\{\bm{g},\bm{e}_{t}\}.$
Soft-MSA then processes $\bm{z}$ via masked projections:
$\bm{q}=(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q})\bm{z},\qquad\bm{k}=(\bm{\theta}^{K}\odot\bm{m}_{t}^{K})\bm{z},$
yielding the attention matrix
$\bm{a}=\mathrm{softmax}\left(\frac{\bm{q}\bm{k}^{\top}}{\sqrt{D/s}}\right).$ (2)
G-prompts impart globally relevant cues, while E-prompts and soft-masks enable fine-grained, task-specific adjustments. This combination yields strong task separation without requiring explicit task identifiers during evaluation.
4.3 Optimization over Frozen Backbones
For task $t$, the predictive model is
$f_{t}(\bm{x})=f_{\phi}\big(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(x)\big),$
where $\phi$ is a lightweight task classifier. Training optimizes:
$\min_{\bm{g},\,\bm{e}_{t},\,\bm{m}_{t},\,\bm{k}_{t},\,\phi}\;\mathcal{L}_{\mathrm{CE}}\big(f_{\phi}(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})),y\big)+\lambda\,\gamma\big(q(\bm{x}),\bm{k}_{t}\big),$ (3)
where $\bm{k}_{t}$ is a task-level key vector and $\gamma$ is a cosine-similarity regularizer, defined as $1-\cos(\cdot,\cdot)$, which facilitates task inference at test time. Only the prompts, soft-masks, and classifier parameters are updated; the backbone remains frozen, yielding low memory cost and consistent forward-time efficiency. The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm˜1 and Algorithm˜2.
Algorithm 1 Training Procedure of DualPrompt-Soft-TF
1:  Input: Pre-trained transformer backbone $f$, classifier $f_{\phi}$;
2:   number of tasks $\mathcal{T}$; training sets $\{\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{n_{t}}\}_{t=1}^{\mathcal{T}}$;
3:   G-Prompt $\bm{g}$, E-Prompts $\bm{E}=\{\bm{e}_{t}\}_{t=1}^{\mathcal{T}}$;
4:   task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$;
5:   soft-networks $\bm{M}=\{\bm{m}_{t}\}_{t=1}^{\mathcal{T}}$;
6:   prompt insertion layers $l^{start}_{g},l^{end}_{g}]$, $l^{start}_{e},l^{end}_{e}]$;
7:   prompting function $f_{\bm{\theta}\odot\bm{m}}$; epochs $\mathcal{K}_{t}$ for task $t$.
8:  Initialize: $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$
9:  for task $t=1,\dots,\mathcal{T}$ do
10:  Select task-specific $\bm{e}_{t}$, soft-network $\bm{m}_{t}$, and key $\bm{k}_{t}$
11:  Construct prompted backbone $f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}$ by attaching:
12:   G-Prompt to layers $l^{start}_{g},l^{end}_{g}]$,
13:   E-Prompt to soft-MSA layers $l^{start}_{e},l^{end}_{e}]$ via $f_{\bm{\theta}\odot\bm{m}_{t}}$
14:  for epoch $s=1,\dots,\mathcal{K}_{t}$ do
15:   Sample mini-batch $B=\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{l}$
16:   for each $(\bm{x},y)$ in $B$ do
17:    Compute prompted representation:
18:     $\bm{h}=f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})$
19:    Compute prediction $\hat{y}=f_{\phi}(\bm{h})$
20:    Compute per-sample loss:
21:     $\mathcal{L}_{x}=\mathcal{L}_{CE}(\hat{y},y)+\lambda\,\gamma(q(\bm{x}),\bm{k}_{t})$
22:   end for
23:   Update $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$ via backpropagation
24:  end for
25:  end for
4.4 Stability and Bounded Parameter Drift
Define the effective parameters for task $t$ as
$\bm{w}=\bm{\theta}\odot\bm{m}_{t},$
so that $\bm{w}^{(1)}=\bm{\theta}$ under mask initialization $\bm{m}_{t}^{(1)}=\mathbf{1}$. A single gradient update yields:
$\bm{w}^{(k+1)}=\bm{w}^{(k)}-\eta\big(\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\big).$
Thus, the optimization path is governed by a rescaled gradient whose magnitude is bounded by the scale of $\bm{\theta}$. Under standard convex–Lipschitz conditions, we obtain the cumulative suboptimality bound:
$\begin{aligned} \sum_{k=1}^{T}&\langle\bm{w}^{(k)}-\bm{w}^{\star},\,\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\rangle\leq\\ &\mathcal{O}\!\left(\frac{\|\bm{w}^{(1)}-\bm{w}^{\star}\|^{2}}{\eta}+\eta\sum_{k=1}^{T}\|\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\|^{2}\right).\end{aligned}$
Table 1: Performances of Vision Class Incremental Learning in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note that Soft-TF performed without FeedForward ($F$) networks. $\ast$ denotes our reproduced results and L[10,11,12] represents Soft-TransFormer of 10, 11, 12 Layers.
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
+ PGP 0.03M Prompt ID 84.34 / 5.59 12.30K / 75 76.12 / 13.26 12.70K / 76 61.40 / 8.03 13.07K / 46
+ Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
+ PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
+ Adapter 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
+ LoRA 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.2 15.09K / 53
+ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
+ Soft-TF-L[10,11,12] (SOTA) 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
Joint of DualPrompt 0.13M - 90.85 - 90.85 - 79.13 -
Joint of DualPrompt+Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
Joint of NoRGa+Soft-TF 6.93M - 98.20 - 98.20 - 90.73 -
Implications for continual learning. The initial distance $\|\bm{w}^{(1)}-\bm{w}^{\ast}\|$ is naturally small across most tasks because the pre-trained backbone already provides strong feature representations. Consequently, by standard convergence guarantees of first-order optimization for Lipschitz-continuous objectives without assuming strong convexity, the suboptimality decreases at a rate of $\mathcal{O}(1/\sqrt{T})$. More importantly, since optimization occurs only in the low-dimensional soft-parameter space $(\bm{m}_{t}^{Q},\bm{m}_{t}^{K},\bm{m}_{t}^{V})$ while $\bm{\theta}$ remains fixed, the entire trajectory $\bm{w}^{(k)}$} stays confined to a narrow region around the pretrained solution manifold. This imposes an inherent bound on parameter drift and explains why Soft-TF reliably preserves global knowledge encoded in $\bm{\theta}$ while still enabling targeted adaptation to each new task. Taken together, these theoretical results (see Section˜A.1) provide a principled foundation for our design choice: initializing masks at $\mathbf{1}$ and fine-tuning only $\bm{m}_{t}$ over a frozen backbone yields stable optimization, bounded deviation from pre-trained representations, and substantially reduced catastrophic forgetting in continual learning.
5 Experiments
We validate our Soft-TF on several benchmark datasets against continuous learning baselines in Class-Incremental Learning (CIL).
5.1 Experimental Settings
Datasets.
We evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks. 2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks. 3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks. To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200.
Baselines. To validate the powerfulness of our method, we compare our results with various CIL baselines including, L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), DualPrompt-PGP (Qiao et al., 2024), and NoRGa (Le et al., 2024). At inference time, the system determines the task identity for an arbitrary test sample $\bm{x}$ to select the appropriate task-specific nuances and ultimately achieve full fine-tuning performance. This is accomplished by using the Prompt ID mechanism, which simply chooses the best-matched task index $t$ by finding the minimum value of a function $\gamma$ comparing the query representation of the sample, $q(\bm{x})$, against all learned task keys $\bm{k}_{t}$, expressed as $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
Table 2: Performances of DualPrompt+Soft-TF in terms of ViTs pretained-dataset (ImageNet-21K) on 10-Split-ImageNet-R.
DualPrompt 10-Split-ImageNet-R
+Soft-TF #Tr.Prams ACC($\uparrow$) Forget($\downarrow$) Tr./Test[sec]
ViT-Ti/16 (5.7M) 0.41M 49.34 9.02 15.10K / 50
ViT-S/16 (22M) 1.71M 65.12 6.67 15.22K / 51
ViT-B/16 (86M) 6.93M 76.62 5.30 15.35K / 52
ViT-L/16 (307M) 12.50M 78.86 4.46 28.57K / 99
Refer to caption Refer to caption
(a) Comparisons with baselines (b) Layer-wise Performances
Figure 4: Layer-wise(L[$\ast$]) Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-Transformer of 10, 11, 12 Layers.
5.2 Performances
Performances of Soft-TF. We compare our Soft-TF with state-of-the-art CIL baselines, as shown in Table˜1. Soft-TF consistently outperforms prompt-based and parameter-efficient fine-tuning methods—including L2P, DualPrompt, NoRGa, Adapter, LoRA, and their corresponding upper-bounds—across accuracy, forgetting, and efficiency metrics. In particular, Soft-TF exhibits substantial improvements over NoRGa, a strong prompt-based baseline that mitigates task interference, yet remains limited by its reliance on global or head-level prompt modulation. By contrast, Soft-TF directly modulates attention projection matrices through task-specific soft subnetworks, enabling fine-grained and continuous adaptation within a frozen Transformer backbone, which results in higher accuracy and significantly reduced forgetting across all benchmarks. Despite using comparable trainable parameter budgets, Soft-TF achieves these gains without additional training or inference cost. Remarkably, Soft-TF even surpasses the Upper-Bound (Joint) of DualPrompt using prompt baselines: DualPrompt and NoRGa on all evaluated benchmarks, as it explicitly identifies and optimizes distinct task-specific soft subnetworks for each incremental task, achieving near-optimal per-task specialization, whereas joint training is constrained to a single shared solution that must compromise across tasks.
Comparisions with baselines. Figure˜4 of (a) shows that Soft-TF consistently outperforms established parameter-efficient tuning baselines - including PGP, Adapter, and LoRA - across all sessions in 10-Split CIFAR-100, demonstrating superior resistance to task interference and better retention of previously learned knowledge. As shown in Figure˜4 of (b) the layer-wise analysis, even when fine-tuning only selected layers (e.g., layers 10, 11, 12), Soft-TF exhibits a range of performance behaviors depending on the chosen layer configuration, yet it still maintains a clear advantage over LoRA in every setting. This indicates that Soft-TF not only provides stronger overall continual learning stability but also preserves its effectiveness under various layer-wise adaptation strategies, highlighting its robustness and versatility.
5.3 Variants of ViTs.
Across different ViT backbones, Soft-TF exhibits consistent advantages in both parameter efficiency and continual learning stability on 10-Split-ImageNet-R, as shown in Table˜2. The lightweight ViT-Ti/16 uses only 0.41M trainable parameters but yields lower accuracy (49.34%) and higher forgetting (9.02), whereas scaling up to ViT-S/16, ViT-B/16, and ViT-L/16 increases trainable parameters to 1.71M, 6.93M, and 12.50M, respectively, resulting in markedly improved performance. In particular, Soft-TF with ViT-L/16 achieves the best overall results - 78.86% accuracy and only 4.46 forgetting - while maintaining comparable training and test times ($\approx$28K/99 sec). Figure˜5 shows that larger backbones maintain higher average accuracy across all 10 sessions, with ViT-L/16 degrading the slowest and sustaining a clear performance margin over ViT-B/16, ViT-S/16, and ViT-Ti/16. Overall performance differences are further influenced by each backbone’s representational capacity and initialization quality, which shape how effectively Soft-TF can adapt while resisting catastrophic forgetting. Moreover, while increasing the number of trainable parameters generally leads to better continual learning performance, this improvement naturally comes with higher computational cost, reflecting the typical trade-off between model capacity and efficiency.
Figure 5: Performances of DualPrompt+Soft-TF: in terms of ViT models on 10-Split-ImageNet-R. The performance of Soft-TF-ViT-B/16 is comparable to that of Soft-TF-ViT-L/16.
5.4 Well-initialized Soft-TF
Random initialization. Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs). The optimal training point is the parameters of a well-pretrained model. Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively. To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table˜3. The results demonstrate that the same well-initialization point leads to independent optimal task performance. Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
Well-initialized LTH (WLTH). Based on Table˜4, we evaluated our Soft-TF on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three distinct pre-trained ViT models (ImageNet-21K, SAM, and DINO) to demonstrate its efficacy, including on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones. For the 10-Split-CIFAR100 benchmark, the ImageNet-21K backbone achieved the highest accuracy (92.35%) and the lowest forgetting rate (2.98%), closely followed by the SAM backbone (92.30% ACC, 2.99% Forget). Similarly, for the 5-Split-CUB200 benchmark, the ImageNet-21K backbone led with 76.17% ACC and 9.04% Forget, which was also nearly matched by the SAM backbone (76.11% ACC, 9.01% Forget). In contrast, the DINO backbone consistently showed the lowest performance across both datasets. These trends are consistent with our theoretical analysis, which predicts faster convergence and bounded parameter drift when optimization is initialized near a well-pretrained solution and confined to a low-dimensional soft-parameter space. These results confirm the effectiveness of Soft-TF and suggest that using well-initialized weights provides superior generalization and performance in continual learning settings.
Table 3: Random initialized Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note DualPrompt’s parameters are pretrained on ImageNet-21K.
10-Split-CIFAR100
Random Init. ACC($\uparrow$) Forget($\downarrow$)
Xavier 90.59 3.85
Kaiming 90.72 3.63
Normal 90.45 3.78
Uniform(1.0, 1.0) 92.35 2.98
Table 4: Performances of Soft-TF in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) on 10-Split-CIFAR100 and 5-Split-CUB200.
DualPrompt 10-Split-CIFAR100 5-Split-CUB200
+Soft-TF ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
ImageNet-21K 92.35 2.98 76.17 9.04
SAM 92.30 2.99 76.11 9.01
DINO 84.50 12.27 69.79 10.93
5.5 Soft-MSA Layers
The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure˜6: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3). In Figure˜6(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1. There is a slight variation between Task 2 and Task 3. However, the overall range remains consistent. $\bm{m}^{QKV}$ acts as a soft gating mechanism, controlling which pre-trained parameters are retained or modified during fine-tuning. In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{\theta}^{QKV}$) is centered around 0, with a Gaussian-like spread. The distribution of soft-masked fine-tuned weight ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{\theta}^{QKV}$ but slightly narrower. The overlap between the two distributions indicates that the fine-tuning process maintains the general structure of the pre-trained weights while introducing task-specific modifications. By using soft-masks $\bm{m}^{QKV}$, Soft-TF ensures task-specific fine-tuning, reducing the risk of catastrophic forgetting and improving performance in various sequential tasks. The task-specific Soft-TF fine-tuning enables the models to infer task ID ideally in following Vision CIL scenarios.
Refer to caption Refer to caption
(a) $\bm{m}^{QKV}$ (b) $\bm{\theta}^{QKV}$ v.s. $\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$
Figure 6: Attention Layer-(L[12]) Histogram Density Estimates of Pre-trained weight ($\bm{\theta}$) and Soft-TF ($\bm{m}$) on 10-Split-CIFAR100.
Refer to caption Refer to caption
(a) DualPrompt-L[12] (b) Soft-TF-L[12]
Figure 7: Attention map comparisons of DualPrompt and Soft-TF: at Layer-(L[12]) on a single category 7 sample $\bm{x}_{i}^{t}$ of 10-Split-CIFAR100.
Attention Layer. We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure˜7. While DualPrompt distributes its attention more uniformly, Soft-TF prioritizes specific patches of the input image, likely those more relevant to Category 7. These attention maps highlight the differences in how DualPrompt and Soft-TF allocate attention in a Vision Transformer. Soft-TF’s sharper focus on essential features aligns with the reported superior performance on the 10-Split-CIFAR100 dataset. This visualization emphasizes the interpretability and task-specific optimization of Soft-TF over the more generalized behavior of DualPrompt.
6 Conclusion
Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformers (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer. Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention. These masks enable smooth and stable task adaptation while preserving shared representations. Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF). Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
Appendix A Appendix
A.1 Analysis of Soft-TransFormers (Soft-TF)
Analysis of Soft-TransFormers for Convex-Lipschitz Functions. To analyze the convergence rate of the Soft-TransFormers (Soft-TF), we limit ourselves to the case of convex-Lipshitz functions along with the analysis (Shalev-Shwartz and Ben-David, 2014). Let $\bm{w}^{\ast}=\{\bm{g}^{\ast},\bm{e}_{t}^{\ast},\bm{m}_{t}^{\ast}\}$ be any vector or an optimal solution and let $B$ be an upper bound on $\bm{w}^{\ast}||$ when $\bm{w}^{(1)}=\bm{0}$. It is convenient to think of $\bm{w}^{\ast}$ as the minimizer of $f(\bm{w})$, but the analysis that follows holds for every $\bm{w}^{\ast}$.
We obtain an upper bound on the sub-optimality of our solution with respect to $\bm{w}^{\ast}$, namely, $f(\bar{\bm{w}})-f(\bm{w}^{\ast})$, where $\bar{\bm{w}}=\frac{1}{T}\bm{w}^{(t)}$. From the definition of $\bar{\bm{w}}$, and using Jensen’s inequality, we have that
$\begin{aligned} f(\bar{\bm{w}})-f(\bm{w}^{\ast})&=f\left(\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{(t)}\right)-f(\bm{w}^{\ast})\\ &\leq\frac{1}{T}\sum_{t=1}^{T}\left(f(\bm{w}^{(t)})\right)-f(\bm{w}^{\ast})\\ &=\frac{1}{T}\sum_{t=1}^{T}\left(f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\right).\end{aligned}$
For every $t$, because of the convexity of $f$, we have that
$f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\leq\left<\bm{w}^{(t)}-\bm{w}^{\ast},\nabla f(\bm{w}^{(t)})\right$>
Combining the preceeding we obtain
$f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\leq\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\nabla f(\bm{w}^{(t)})\right$>
To bound the right-hand side we rely on the following lemma:
Lemma A.1.
Let $\bm{v}_{1},\cdots,\bm{v}_{T}$ be an arbitrary sequence of vectors. Any algorithm with an well initialization (pre-trained model) $\bm{w}^{(1)}\neq\bm{0}$ and an update rule of the form
$\bm{w}^{(t+1)}=\bm{w}^{(t)}-\eta\bm{v}_{t}$
[…]
where $\bm{w}_{m}\neq\bm{w}_{q}$ since $\bm{m}$ is learnable parameters in Soft-TransFormers. Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of LABEL:eq:lr_params.
Theorem A.2.
[…]
$\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
Refer to caption Refer to caption
(a) L2P-PGP v.s. L2P+Soft-TF (b) DualPrompt v.s. DualPrompt+Soft-TF
Figure 8: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-TransFormer of 10, 11, 12 Layers.
Proof.
Using algebraic manipulations (completing the square), we obtain:
[…]
where the last equality follows from the definition of the update rule. Summing the equality over $t$, we have
[…]
The first sum on the right-hand side is a telescopic sum that collapses to
[…]
Plugging this in Equation, we have
[…]
where the last equality is due to the definition $\bm{w}^{(1)}=\bm{0}$. This proves the first part of the lemma. The second part follows by upper bounding $\bm{w}||$ by $B$, $\bm{v}_{t}||$ by $\rho$, deciding by $T$, and plugging in the value of $\eta$.
Table 5: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R. Exemplar means the total buffer size for rehearsal methods.
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
BiC 5,000 - 81.42 17.31 73.02 6.23 64.63 22.25
DER++ 5,000 - 83.94 14.55 - - 66.73 20.67
iCaRL 5,000 - 66.00 5.33 78.02 5.80 - -
DER+MCG 2,000 - 67.62 14.64 65.84 13.72 - -
BiC 1,000 - 66.11 35.24 63.12 21.89 52.14 36.70
DER++ 1,000 - 61.06 39.87 - - 55.47 34.64
iCaRL 1,000 - 61.25 14.19 71.32 15.98 - -
FT - - 33.61 86.87 33.52 53.69 28.87 63.80
EWC - - 47.01 33.27 36.73 35.19 35.00 56.16
LWF - - 60.69 27.77 39.12 57.91 38.54 52.37
DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
+ PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
+ Soft-TF-L[3,4,5] 6.93M Prompt ID 91.77 / 3.37 13.87K / 80 94.43 / 2.02 15.60K / 104 74.70 / 6.46 15.35K / 52
+ Soft-TF-L[3,4,5] 6.93M Gradient ID 93.76 / 1.83 13.87K / 130 95.38 / 1.73 15.60K / 163 82.15 / 2.20 15.35K / 82
+ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
+ Soft-TF-L[10,11,12] 6.93M Gradient ID 97.87 / 0.21 13.87K / 130 99.05 / 0.24 15.60K / 163 82.38 / 0.59 15.35K / 82
Upper-Bound of DualPrompt 0.03M - 90.85 - 90.85 - 79.13 -
Upper-Bound of Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$, we have
[…]
where $\bm{v}_{t}$ is an arbitrary $t$-th vector and $\bm{w}^{(1)}\neq\bm{0}$ since $\bm{w}^{Q}$ and $\bm{q}^{K}$ are pre-trained parameters.
however, in term of prompt $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$, we have
[…]
where $\bm{v}_{t}$ is an arbitrary $t$-th vector of prompt and $\bm{w}^{(1)}\neq\bm{0}$ since $(\bm{w}\odot\bm{1})^{Q}$ and $(\bm{w}\odot\bm{1})^{K}$ are pre-trained parameters, $\bm{w}^{Q}$ and $\bm{w}^{K}$, respectively.
[…]
$\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
A.2 Experimental Details
For fair comparisons with the baselines (Wang et al., 2022d, c; Qiao et al., 2024), we use ViT B/16 (Dosovitskiy et al., 2021) pre-trained on ImageNet-21K as our image encoder, which is kept frozen during training. We train and test on a single Quadro RTX 8000-48GB GPU for baselines and our Soft-TransFormers with Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
We adhere to the experimental settings outlined by (Qiao et al., 2024) to validate our method’s effectiveness. When comparing our approach with L2P-PGP and Soft-Transformer on the 10/20-Split-CIFAR100 and 10-Split-TinyImageNet datasets, we train the network for 5 epochs with a batch size of 16 and set the prompt length to 5. For the 10-Split-ImageNet-R dataset, we use 50 epochs, a batch size of 16, and a prompt length of 30. In comparison with DualPrompt-PGP and Soft-TransFormers on the 10/20-Split-CIFAR100 dataset, we train the network for 20 epochs with a batch size of 24 and set the expert prompt length to 5. For the 10-Split-TinyImageNet dataset, we use 5 epochs, a batch size of 24, and an expert prompt length of 5. For the 10-Split-ImageNet-R dataset, we set the epochs to 50, the batch size to 24, and the expert prompt length to 20. Additionally, in all benchmark data sets, the general prompt length is set to 5, and the location inserted into the prompt is kept consistent.
Table 6: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R.
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
+ Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
+ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
S-Prompt 0.03M Prompt ID 87.57 / 3.63 12.12K / 76 84.90 / 7.05 11.60K / 78 74.25 / 4.73 13.10K / 47
CODA-Prompt 0.03M Prompt ID 86.94 / 4.04 12.12K / 76 84.70 / 7.11 11.60K / 78 74.26 / 5.17 13.10K / 47
HiDe-Prompt 0.03M Prompt ID 92.61 / 1.50 12.12K / 76 94.28 / 1.78 11.60K / 78 75.06 / 4.09 13.10K / 47
NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
+ Soft-TF-L[10,11,12] 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
Comparisons with Prompt-based Baselines. We evaluate the performance of Soft-TF against representative prompt-based continual learning methods, including L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), CODA-Prompt (Smith et al., 2023b), S-Prompt (Wang et al., 2022a), HiDe-Prompt (Wang et al., 2023), and NoRGa (Le et al., 2024), as summarized in Table˜6. While these approaches demonstrate that improved prompt representations generally lead to better continual learning performance, their adaptation capacity remains fundamentally constrained by prompt-level conditioning. In contrast, Soft-TF consistently achieves state-of-the-art results by directly modulating attention projections through task-specific soft subnetworks, providing substantially greater representational expressiveness than prompts alone. These results indicate that, even when compared to strong and diverse prompt-based baselines, Soft-TF offers a more effective and scalable mechanism for continual adaptation.
Task Inference. At the inference time, we infer task identity for arbitrary pieces of task samples $\bm{x}$ for finding the proper task nuances and demonstrating full fine-tuning results. We summarize the following two methods:
  • Prompt ID: For a test example $\bm{x}$, we simply choose the best matched task index via $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
[…]
Using the Gradient ID inference method, as shown in Table˜5, Soft-TF demonstrates superior average performance and minimal forgetting compared to the Prompt ID approach, even with a slight increase in computational load at inference time. This method is not only valuable for applications requiring the highest level of performance but also serves as a foundational study for future research into prompt learning and task-ID inference.
Layer-wise Inspections. We analyze the layer-wise performance of Soft-Transformer with respect to L2P and DualPrompt on the 10-Split-CIFAR100 dataset to identify the optimal configurations, as shown in Figure˜8. Our observations reveal that the global prompt in DualPrompt influences Soft-Transformer’s performance differently in L2P and DualPrompt settings. In L2P-PGP, the best performance was achieved with Soft-TransFormers applied to the lower layers ((a) L2P-Soft-TF-L[1,2]-PGP), whereas in DualPrompt, the higher layers ((b) DualPrompt-Soft-TF-L[10,11,12]) yielded the best results. Notably, DualPrompt-Soft-TF-L[10,11,12] without PGP demonstrated impressive performance, achieving almost zero forgetting (0.21). These findings suggest that our approach could significantly enhance the effectiveness of large-scale Transformer models in continual learning scenarios.
Peudo Codes. The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm˜1 and Algorithm˜2. We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$. The input $\bm{x}$ from the $t$-th task is transformed using $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$ and then passed to the classification head $f_{\phi}$, parameterized by $\phi$, for prediction. Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner.
Refer to caption Refer to caption
(a) Attn.QKV ($\bm{m}^{QKV}$) (b) Attn.Proj ($\bm{m}^{O}$)
Refer to caption Refer to caption
(c) Mlp.fc1 ($\bm{m}^{fc1}$) (d) Mlp.fc2 ($\bm{m}^{fc2}$)
Figure 9: Layer-(L[12]) Histogram Density Estimates of DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
Refer to caption Refer to caption
(a) $\bm{\theta}^{QKV}_{Attn}$ v.s. $\bm{\theta}^{QKV}_{Attn}\odot\bm{m}^{QKV}_{Attn}$ (b) $\bm{\theta}^{O}_{Proj}$ v.s. $\bm{\theta}^{O}_{Proj}\odot\bm{m}^{O}_{Proj}$
Refer to caption Refer to caption
(c) $\bm{\theta}^{fc1}$ v.s. $\bm{\theta}^{fc1}\odot\bm{m}^{fc1}$ (d) $\bm{\theta}^{fc2}$ v.s $\bm{\theta}^{fc2}\odot\bm{m}^{fc2}$
Figure 10: Layer-(L[12]) Histogram Density Estimates of Pre-trained Weight and DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
Density of Parameters. We inspect the histogram density estimate of the last (12) layer’s parameters of DualPrompt-Soft-TF: attention of QKV ((a) Attn.QKV) and Projection ((b) Attn.Proj) and multi-layer perception (Mlp) of fc1 and fc2, as shown in Figure˜9. Attn’s QKV parameters have the largest variance among the parameter densities, while MLP-FC2’s are the smallest. From this observation, we conclude that fine-tuning Attn’s QKV is required to achieve optimal task performance. In other words, QKV’s parameters are more critical than others.
Pre-trained Parameters vs. Soft-TF. We inspect the histogram density estimate of the last (12) layer’s parameters of pre-trained model and DualPrompt-Soft-TF: attention of QKV ((a) Attn.QKV) and Projection ((b) Attn.Proj) and multi-layer perception (Mlp) of fc1 and fc2, as shown in Figure˜10. The most parameters of DualPrompt-Soft-TF are trained around zero-values. Particularly, the difference between pre-trained model’s parameters and Soft-TF is distinctive at QKV module.
Additional Comparisons with Prompt-based Baselines. As shown in Table˜7, Soft-Transformers (Soft-TF) achieve a substantial performance margin over established prompt-based continual learning methods on 10-Split-CIFAR100. Whereas conventional prompt learners such as L2P, DualPrompt, ESN, CODA-Prompt, and CPrompt obtain accuracies in the range of 77–87% with forgetting between 5–7%, Soft-TF reaches 92.35% accuracy and reduces forgetting to only 2.98. This improvement is not incremental but a clear step-change: Soft-TF reduces forgetting by nearly half while surpassing the best prompt-based accuracy by more than 5 percentage points. These results highlight a fundamental limitation of prompt-only mechanisms, which rely solely on input-level conditioning and thus struggle to capture fine-grained, task-specific distinctions under heterogeneous task streams. By contrast, Soft-TF leverages task-specific soft subnetworks that directly modulate attention projections, offering richer representational flexibility while maintaining the frozen backbone. The strong gains over diverse prompt baselines underscore that Soft-TF provides a more expressive and stable adaptation mechanism than prompting alone, even when evaluated under random initialization.
Algorithm 2 DualPrompt-Soft-TF at Test Time
1:  Given: Pre-trained $f$; trained task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$;
2:   learned soft-networks $\bm{M}=\{\bm{m}_{t}\}_{t=1}^{\mathcal{T}}$;
3:   prompt insertion layers $l^{start}_{g},l^{end}_{g}]$, $l^{start}_{e},l^{end}_{e}]$;
4:   prompting function $f_{\bm{\theta}\odot\bm{m}}$.
5:  Input: Test example $\bm{x}$ (from mini-batch $\bm{b}$)
6:  Task inference method:
7:   Choose either (1) Prompt ID or (2) Gradient ID
8:  (1) Prompt ID
9:   Compute query feature $q(\bm{x})$
10:   Select task index:
11:   $t_{\bm{x}}=\arg\min_{t}\gamma\big(q(\bm{x}),\bm{k}_{t}\big)$
12:  (2) Gradient ID
13:   Initialize task weights $\{\alpha_{t}\}_{t=1}^{\mathcal{T}}$,
14:   with $\alpha_{t}=1/\mathcal{T}$ and $\sum_{t}\alpha_{t}=1$
15:   Compute soft-merged network:
16:   $\bm{m}_{\alpha}=\sum_{t=1}^{\mathcal{T}}\alpha_{t}\bm{m}_{t}$
17:   Compute loss on $\bm{x}$:
18:   $\mathcal{L}=\mathcal{H}\big(f_{\bm{\theta}\odot\bm{m}_{\alpha}}(\bm{x})\big)$
19:   Select task index via gradient signal:
20:   $t_{\bm{x}}=\arg\min_{t}\frac{\partial\mathcal{H}}{\partial\alpha_{t}}$
21:  Prompt selection
22:  Select E-Prompt $\bm{e}_{t_{\bm{x}}}$ and soft-network $\bm{m}_{t_{\bm{x}}}$
23:  Construct prompted model $f_{\bm{g},\bm{e}_{t_{\bm{x}}},\bm{m}_{t_{\bm{x}}}}$:
24:   Attach $\bm{g}$ at layers $l^{start}_{g},l^{end}_{g}]$
25:   Attach $\bm{e}_{t_{\bm{x}}}$ at layers $l^{start}_{e},l^{end}_{e}]$
26:   Apply soft-transformer routing via $f_{\bm{\theta}\odot\bm{m}_{t_{\bm{x}}}}$
27:  Prediction:
28:   $\hat{y}=f_{\bm{g},\bm{e}_{t_{\bm{x}}},\bm{m}_{t_{\bm{x}}}}(\bm{x})$
Table 7: Comparisons with others on 10-Split-CIFAR100.
10-Split-CIFAR100
Method ACC($\uparrow$) Forget($\downarrow$)
L2P 83.38 5.88
DualPrompt 86.61 5.86
ESN 86.42 6.08
CODA-Prompt 85.73 7.13
CPrompt 77.14 5.97
DualPrompt+Soft-TF 92.35 2.98
Additional Results on CUB-200 & 5-Datasets. As shown in Table˜8, combining NoRGa with Soft-TF yields consistently superior fine-tuning performance across both Split CUB-200 and 5-Datasets under different pretraining regimes. While NoRGa already demonstrates strong continual adaptation by mitigating task interference, the integration of Soft-TF further unlocks the expressive capacity of the frozen Transformer backbone by directly modulating attention projections through task-specific soft subnetworks. This results in a substantial and consistent accuracy gain over NoRGa alone, achieving near-ceiling performance on all benchmarks (e.g., 97.90% on Split CUB-200 and 95.68% on 5-Datasets with Sup-21K pretraining). The improvements are not marginal but systematic, indicating that Soft-TF provides a powerful complementary mechanism to NoRGa by enabling fine-grained, task-aligned parameter adaptation beyond prompt- or gradient-level regularization. These results clearly demonstrate that NoRGa+Soft-TF delivers an overwhelmingly strong fine-tuning strategy, effectively bridging robust continual learning with highly expressive task-specific specialization.
Table 8: Final average accuracy (FA) on Split CUB-200 and 5-Datasets.
Method Split CUB-200 5-Datasets
Sup-21K iBOT-21K Sup-21K iBOT-21K
L2P 75.46 46.60 81.84 82.25
DualPrompt 77.56 45.93 77.91 68.03
S-Prompt 77.13 44.22 86.06 77.20
CODA-Prompt 74.34 47.79 64.18 51.65
HiDe-Prompt 86.56 78.23 93.83 94.88
NoRGa 90.90 80.69 94.16 94.92
NoRGa + Soft-TF 97.90 90.82 95.68 95.42